# Multi-Persona Thinking for Bias Mitigation in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) exhibit significant social biases that can perpetuate harmful stereotypes and unfair outcomes. In this paper, we propose Multi-Persona Thinking (MPT), a novel inference-time framework that leverages dialectical reasoning from multiple perspectives to reduce bias. MPT guides models to adopt contrasting social identities (e.g., male and female) along with a neutral viewpoint, and then engages these personas iteratively to expose and correct biases. Through a dialectical reasoning process, the framework transforms the potential weakness of persona assignment into a strength for bias mitigation. We evaluate MPT on two widely used bias benchmarks across both open-source and closed-source models of varying scales. Our results demonstrate substantial improvements over existing prompting-based strategies: MPT achieves the lowest bias while maintaining core reasoning ability.
@@
Yuxing Chen    Guoqing Luo    Zijun Wu    Lili Mou* Dept. Computing Science & Alberta Machine Intelligence Institute, University of Alberta *Canada CIFAR AI Chair {yuxing2, gluo, zijun4, lmou}@ualberta.ca
@@ 1 Introduction
Large Language Models (LLMs) have shown strong performance across many natural language processing tasks, including text generation, question answering, and dialogue systems (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023). However, researchers have shown that LLMs also exhibit biases, such as misrepresentations and stereotypes (Bender et al., 2021; Hutchinson et al., 2020). Such biases can be either explicit, involving overt associations, or implicit, reflecting unconscious patterns that influence reasoning and behavior (Greenwald et al., 1998; Brownstein, 2025). When embedded in model outputs, such biases reinforce stereotypes and may cause unfair or harmful results when applied in sensitive domains such as hiring, education, and healthcare (Raghavan et al., 2020; Selwyn, 2019; Davenport and Kalakota, 2019).
@@
Bias mitigation techniques exist in different stages of the LLM workflow, including pre-processing of training data (Lu et al., 2020; Garimella et al., 2022), in-training methods modifying model architectures or objectives (Lauscher et al., 2021; Ouyang et al., 2022), inference-time interventions like self-debiasing prompts (Schick et al., 2021), and post-processing filters that rewrite outputs (Tokpo and Calders, 2022; He et al., 2021).
@@
Among these categories, prompt-based debiasing has received considerable attention because it is computationally efficient and does not require access to model parameters. For example, self-debiasing prompts encourage the model to first reflect on potential harmful stereotypes and then adjust its responses accordingly (Schick et al., 2021; Gallegos et al., 2025; Ganguli et al., 2023; Zhao et al., 2025). Zheng et al. (2024) and Gupta et al. (2024) explore persona-assigning and role-playing prompting for enhancing reasoning ability, where LLMs are instructed to respond based on a certain social identity or domain expert. However, they only assign one persona/role in the prompt, so the LLM may be biased toward the assigned persona/role.
@@
In this paper, we propose a Multi-Persona Thinking (MPT) approach, an inference-time framework for mitigating social bias in LLMs. Specifically, we guide an LLM to reason from the perspectives of different personas that are related to the bias category (e.g., male and female). Inspired by multi-agent debate (Du et al., 2023), we let different personas exchange, compare, and refine their viewpoints. This process can gradually reveal and reduce biases through the interaction of diverse perspectives. Unlike prior work (Borah and Mihalcea, 2024) that fine-tunes models to alleviate bias, our method applies dialectical reasoning directly at inference time, which makes MPT a lightweight and applicable method.
@@
Our experiments show that our approach largely outperforms previous methods on prompting strategy. On the Bias Benchmark for QA (BBQ; Parrish et al., 2022), MPT achieves the lowest bias scores on ambiguous questions while preserving core reasoning abilities on disambiguated ones. On the StereoSet (Nadeem et al., 2021), MPT improves accuracy by 28% and reduces bias by 36% relative to the next-best methods. Further, we show that MPT can be combined with other techniques to achieve even higher performance in both accuracy and fairness.
@@ 2 Related Work
2.1 Bias Mitigation in Language Models
Social biases in LLMs originate from multiple sources, such as skewed training corpora, annotator subjectivity, and amplification during fine-tuning (Bender et al., 2021; Hutchinson et al., 2020). These biases can be explicit, involving overt stereotypical associations, or implicit, reflecting underlying statistical patterns in the model’s representations (Greenwald et al., 1998; Brownstein, 2025). Recent studies also show that models continue to exhibit such biases even after safety alignment (Ganguli et al., 2023; Salewski et al., 2023), making mitigation an ongoing challenge.
Approaches to bias reduction in LLMs can be categorized by the stage of the model pipeline (Gallegos et al., 2024; Blodgett et al., 2020; Li et al., 2023). Pre-processing methods focus on refining the training data before the model is built, such as filtering biased text and augmenting the dataset with more balanced examples to ensure fair representation (Lu et al., 2020; Garimella et al., 2022). In-training methods target the model’s learning phase. For example, Lauscher et al. (2021) modify the model architecture to isolate and control biases, and Ouyang et al. (2022) adjust the training objectives to penalize unfair results. Inference-time methods are used when the trained model is generating a response. These techniques include altering the predicted token distribution to steer the output away from biased language or carefully tuning the prompt to guide the model toward a less biased response (Chung et al., 2023; Schick et al., 2021; Gallegos et al., 2025; Luo et al., 2025). Finally, post-processing methods act as a filter on the finished output by directly rewriting or removing biased content after it has been generated (Tokpo and Calders, 2022; He et al., 2021).
2.2 Prompting & Debiasing
Prompting-based approaches have emerged as a particularly promising direction for bias mitigation due to their computational efficiency and broad applicability. For example,  Ganguli et al. (2023) and  Gallegos et al. (2025) apply the self-reflection (Madaan et al., 2023) technique for debiasing, where a model is prompted to critique its own reasoning for potential stereotypes before providing a final, revised answer.
Salewski et al. (2023) explore persona-assigning or role-playing prompts to instruct models to respond from the perspective of a domain expert or a specific social identity. They show that, although expert roles can enhance models’ coherence or factual grounding, assigning social identity personas may introduce or amplify associated social biases. Subsequent studies further show that relying on a single, static persona can inadvertently reinforce stereotypes or content-dependent biases, demonstrating the limitation and fragility of single-perspective prompting in debiasing tasks (Zheng et al., 2024; Gupta et al., 2024).
Our work uses multiple diverse personas, which differs from existing single-persona approaches. We are inspired by psychological research on perspective-taking and cognitive debiasing (Galinsky and Moskowitz, 2000; Soll et al., 2015), which suggests that considering situations from diverse viewpoints leads to less biased understandings. Our intuition is that stereotypes and biases from different persona viewpoints can be revealed and removed through an iterative dialectical reasoning. However, MPT is different from existing multi-agent frameworks in its interaction goal and structural design. Unlike standard multi-agent debate Du et al. (2023) which mainly focuses on factual consistency, our framework explicitly assigns contrasting social identities to surface latent stereotypes. Borah and Mihalcea (2024) find that multi-agent interactions tend to amplify gender bias. We attribute this divergence to MPT’s distinct design: it replaces persuasive debate with collaborative bias mitigation, and employs additional neutral persona to anchor objective synthesis. Moreover, while Borah and Mihalcea (2024) rely on fine-tuning to mitigate gender bias, MPT shows its effectiveness across a broader categories of social biases as an inference-time framework.
3 Approach
@@
Table 1: Notations for counts in each case (Jin et al., 2024). Amb and Dis refer to ambiguous and disambiguated setups. B and cB refer to the biased and counter-biased answers, respectively. Unk means unknown. Correct predictions are underlined.
Question: Who was not comfortable using the phone?
[…]
This section introduces our approach in detail. Section 3.1 defines the problem formulation, and evaluation metrics including accuracy and diff-bias scores. Section 3.2 presents our proposed Multi-Persona Thinking (MPT) framework.
@@ 3.1 Problem Formulation
In this work, we adopt the widely used problem formulation for LLM debiasing (Parrish et al., 2022; Nadeem et al., 2021; Shaikh et al., 2023; Yang et al., 2025). Each instance consists of a context $c$ and a question $q$, involving two contrasting groups (target group and counter-target group). Three answer choices $\{a_{0},a_{1},a_{2}\}$ are given, referring to a target-biased option, a counter-target-biased option, and an unknown option. Example 1 in Table 2 describes a scenario related to age stereotypes, where the correct answer should be $a_{2}$ (unknown) as other options cannot be inferred from the context. This is referred to as an ambiguous setup. The context can be disambiguated, where $c$ contains additional information that explicitly supports one of the answers. In Example 2, the answer clearly becomes “The grandson” with the additional context information.
To measure the bias in LLMs, we use the accuracy and diff-bias score (Jin et al., 2024) based on the notations shown in Table 1.
@@ 3.2 Multi-Persona Thinking Framework
We propose the Multi-Persona Thinking (MPT) framework based on the assumption that a single LLM encodes a range of social perspectives and associated biases inherently  Gupta et al. (2024). MPT provides a procedure to allow the internal perspectives to exchange viewpoints and reconcile bias during the model’s own reasoning process. We first instantiate different personas with specific system prompts. Then we instruct the model to reason from these perspectives as a simulated self-debate. Finally, the model reviews these viewpoints and provides a final self-debiased answer. In this way, our framework is lightweight and broadly applicable. Details of our approach are introduced below.
@@ 3.2.1 Persona Initialization
For each input example, we initialize a set of three personas $P=\{p_{1},p_{2},p_{3}\}$: two representing the contrasting social groups identified in the dataset metadata (e.g., male vs. female), and one neutral general public persona. Such personas are designed for creating a controlled conflict Galinsky and Moskowitz (2000): the target and counter-target personas allow the model to reveal potential latent biases from opposite perspectives, whereas the neutral persona ensures that the reasoning process is not dominated by either group. Overall, this design encourages balanced evaluation by explicitly aligning opposing perspectives with a “common ground” neutral viewpoint for the final synthesis.
In our approach, we instantiate each persona $p_{i}$ through a persona-assigning system prompt. The details of the prompts are shown in Appendix C.
@@ 3.2.2 Iterative Reasoning
At the beginning ($t=0$), the model independently adopts each persona $p_{i}$ to generate an initial textual response $r_{i}^{(0)}$ and answer $y_{i}^{(0)}\in\{a_{0},a_{1},a_{2}\}$ based on the question prompt $P_{\text{question}}^{(0)}$:
@@
Following the initial generation, the process proceeds through $R$ rounds of dialectical reasoning: In each subsequent round ($t>0$), the model adopting persona $p_{i}$ is instructed to generate a refined response $r_{i}^{(t)}$ and answer $y_{i}^{(t)}$, given a review prompt $P_{\text{review}}^{(t)}$ containing responses from all personas in the previous round, $\{r_{j}^{(t-1)}\}_{j\in\{1,2,3\}}$:
@@
The initialization aims to capture the instinctive, potentially biased answer from each perspective. This makes the model’s internal stereotypes and bias explicit and observable. The subsequent $R$ rounds simulate a cognitive debiasing (Soll et al., 2015): the model is instructed to engage in a self-correction by reasoning from different personas to review and confront the contrasting arguments.
@@ 3.2.3 Final Aggregation
In the final aggregation step, the model is no longer constrained by explicit identities, and it acts as a neutral “judge”. It evaluates the merits of the final arguments presented by all the personas, and integrates them into the final answer. This encourages the model to select the most logical and fact-based conclusion, rather than simple majority voting, ultimately enhancing both the quality and the objectivity of the output.
@@ 4.1 Datasets
To evaluate the effectiveness of our approach, we adopt two widely used stereotyping benchmarks: BBQ (Parrish et al., 2022) and StereoSet (Nadeem et al., 2021).
@@
StereoSet is a benchmark comprising 16,995 crowdsourced Context Association Tests (CAT) that measures language modeling ability and stereotype bias across four domains: gender, profession, race, and religion. The benchmark includes two task formats: word-level (fill-in-the-blank) and sentence-level (choose the most logical continuation). In its original form, each instance provides a context with three candidate associations: one stereotypical, one counter-stereotypical, and the other unrelated. To make our evaluation focus on bias measurement rather than general language modeling, we adapt the test set by replacing the “unrelated” option with an “unknown”. More details are provided in Appendix A.
@@ 4.2 Baselines
To evaluate the effectiveness of our Multi-Persona Thinking (MPT) framework, we select representative baselines that are categorized by their underlying prompting strategy.
Direct prompting. The model is prompted in a single step without iterative refinement. We consider three variants to test the performance of single-turn instruction: (i) Standard prompting, which provides only the task description. (ii) Explicit debiasing, which adds explicit instructions to avoid bias. (iii) Persona-based, which instructs the model to avoid bias by parallel thinking from the perspectives of the two contrasting social groups.
Self-consistency (Wang et al., 2022). This approach samples multiple independent outputs from direct-prompting under stochastic decoding. The final answer is determined by majority voting. We apply self-consistency to all three variants of direct prompting.
Re-prompting (SC; Gallegos et al., 2025). Re-prompting is a two-stage prompting strategy where the model first produces an initial answer, and then is re-invoked with additional explicit debiasing or persona-based instructions to reconsider and potentially revise its response.
Multi-agent debate (MAD; Du et al., 2023). MAD is a framework where multiple model instances engage in debate and refine their reasoning through iterative exchanges. Different from our method, this baseline does not explicitly model diverse social perspectives for bias mitigation.
To ensure a fair comparison, we use the same standard system prompt for answer generation step in all methods and final aggregation step in MPT. All the prompts are detailed in Appendix C.

### v3

@@ Abstract
Large Language Models (LLMs) exhibit social biases, which can lead to harmful stereotypes and unfair outcomes. We propose Multi-Persona Thinking (MPT), a simple inference-time framework that reduces social bias by encouraging reasoning from multiple perspectives. MPT guides the model to consider contrasting social identities, such as male and female, together with a neutral viewpoint. These viewpoints then interact through an iterative reasoning process to identify and correct biased judgments. This design transforms the potential weakness of persona assignment into a mechanism to mitigate bias. We evaluate MPT on two widely used bias benchmarks with both open-source and closed-source models. Our results show that MPT achieves a lower bias than the existing prompting-based methods while maintaining the core reasoning ability.  [^1] Our code is available at https://github.com/MANGA-UOFA/multi-persona-thinking
@@
Yuxing Chen    Guoqing Luo    Zijun Wu    Lili Mou Dept. Computing Science, Alberta Machine Intelligence Institute (Amii) University of Alberta, Canada Canada CIFAR AI Chair, Amii {yuxing2, gluo, zijun4}@ualberta.cadoublepower.mou@gmail.com
@@ 1 Introduction
Large Language Models (LLMs) have shown strong performance in many natural language processing tasks, including text generation, question answering, and dialogue systems (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023). However, researchers have shown that LLMs also exhibit biases, such as misrepresentations and stereotypes (Hutchinson et al., 2020; Bender et al., 2021; Wu et al., 2025). Such biases can be explicit, involving overt associations, or implicit, reflecting unconscious patterns that influence reasoning and behavior (Greenwald et al., 1998; Brownstein, 2025). When reflected in model outputs, such biases can reinforce stereotypes and cause unfair or harmful results in sensitive domains such as hiring, education, and healthcare (Raghavan et al., 2020; Selwyn, 2019; Davenport and Kalakota, 2019).
@@
Bias mitigation techniques have been developed for different stages of the LLM workflow, including pre-processing of training data (Lu et al., 2020; Garimella et al., 2022), in-training methods modifying model architectures or objectives (Lauscher et al., 2021; Ouyang et al., 2022), inference-time interventions like self-debiasing prompts (Schick et al., 2021), and post-processing filters that rewrite outputs (He et al., 2021; Tokpo and Calders, 2022).
@@
Among these categories, prompt-based debiasing has received considerable attention because it is computationally efficient and does not require access to model parameters. For example, self-debiasing instructions encourage the model to first reflect on possible harmful stereotypes and then adjust its responses accordingly (Schick et al., 2021; Ganguli et al., 2023; Gallegos et al., 2025; Zhao et al., 2025). Gupta et al. (2024) and Zheng et al. (2024) explore persona assignment and role-playing prompts to improve reasoning ability, where LLMs are instructed to respond based on a certain social identity or domain expert. However, these methods assign only a single persona or role in the prompt, which may bias the LLM toward that assigned perspective.
@@
In this paper, we propose Multi-Persona Thinking (MPT), an inference-time framework for mitigating social bias in LLMs. Specifically, we guide an LLM to reason from the perspectives of different personas that are related to the bias category (e.g., male and female). Inspired by the multi-agent debate (Du et al., 2023), we prompt a single LLM to simulate perspectives from different personas, and internally exchange and refine them through self-deliberation. This process can gradually reveal and reduce biases through the interaction of diverse perspectives. Unlike previous work (Borah and Mihalcea, 2024) that fine-tunes models to alleviate bias, our method applies dialectical reasoning directly at inference time, making MPT a lightweight and practical method.
@@
Our experiments show that our approach largely outperforms previous prompting methods. On Bias Benchmark for QA (BBQ; Parrish et al., 2022), MPT achieves the lowest bias scores on ambiguous questions while preserving the core reasoning abilities on disambiguated ones. On StereoSet (Nadeem et al., 2021) with Llama, MPT improves accuracy by 30% and reduces bias by 43% relative to the next-best methods. Furthermore, we show that MPT can be combined with other techniques to achieve a stronger performance in both accuracy and fairness.
@@ 2 Related Work
Question: Who was not comfortable using the phone?
[…]
@@
Table 2: Notation for counts in each case (Jin et al., 2024). Amb and Dis refer to ambiguous and disambiguated setups. B and cB refer to the biased and counter-biased answers, respectively. Unk means unknown. Correct predictions are underlined.
2.1 Bias Mitigation in Language Models
Social biases in LLMs originate from multiple sources, such as skewed training corpora, annotator subjectivity, and amplification during fine-tuning (Hutchinson et al., 2020; Bender et al., 2021). These biases can be explicit, involving overt stereotypical associations, or implicit, reflecting underlying statistical patterns in the model’s representations (Greenwald et al., 1998; Brownstein, 2025). Recent studies also show that models continue to exhibit such biases even after safety alignment (Ganguli et al., 2023; Salewski et al., 2023), making mitigation an ongoing challenge.
Approaches to bias reduction in LLMs can be categorized by stage of the model pipeline (Blodgett et al., 2020; Li et al., 2023; Gallegos et al., 2024). Pre-processing methods focus on refining the training data before the model is built, such as filtering biased text and augmenting the dataset with more balanced examples to ensure fair representation (Lu et al., 2020; Garimella et al., 2022). In-training methods target the learning phase of the model. For example, Lauscher et al. (2021) modify the model architecture to isolate and control biases, and Ouyang et al. (2022) adjust the training objectives to penalize unfair results. Inference-time methods are applied when the trained model is generating a response. These techniques include altering the predicted token distribution to steer the output away from biased language or carefully designing the prompt to guide the model toward a less biased response (Schick et al., 2021; Chung et al., 2023; Gallegos et al., 2025). Finally, post-processing methods act as a filter on the finished output by directly rewriting or removing biased content after it has been generated (He et al., 2021; Tokpo and Calders, 2022).
2.2 Prompting & Debiasing
Prompting based approaches have become a promising direction for bias mitigation due to their computational efficiency and broad applicability. For example, Ganguli et al. (2023) and Gallegos et al. (2025) apply self-reflection (Madaan et al., 2023) to debias, where a model is prompted to critique its own reasoning for potential stereotypes before providing a final revised answer.
Salewski et al. (2023) explore persona-assignment or role-playing prompts that instruct models to respond from the perspective of a domain expert or a specific social identity. They show that, although expert roles can enhance models’ coherence or factual grounding, assigning social identity personas may introduce or amplify associated social biases. Subsequent studies further show that relying on a single static persona can inadvertently reinforce stereotypes or content-dependent biases, demonstrating the limitations and fragility of single-perspective prompting in debiasing tasks (Zheng et al., 2024; Gupta et al., 2024).
Our work uses multiple diverse personas, which differs from existing single-persona approaches. We are inspired by psychological research on perspective-taking (Galinsky and Moskowitz, 2000) and cognitive debiasing (Soll et al., 2015), which suggests that considering situations from diverse viewpoints can lead to a less biased understanding. Our intuition is that stereotypes and biases associated with different personas can be surfaced and reduced through iterative dialectical reasoning.
MPT also differs from existing multi-agent frameworks in both interaction goal and structural design. Unlike the standard multi-agent debate (Du et al., 2023), which mainly aims to improve factual consistency, our framework assigns contrasting social identities explicitly to surface latent stereotypes. This distinction is important in the bias setting. For example, Borah and Mihalcea (2024) find that multi-agent interactions can amplify gender bias. We believe that MPT avoids this problem because it is designed for collaborative bias mitigation rather than persuasive debate. It also includes a neutral persona to support a more objective final synthesis. Furthermore, while Borah and Mihalcea (2024) use fine-tuning to mitigate gender bias, MPT is an inference-time framework that generalizes to a broader range of social biases.
3 Approach
This section introduces our approach in detail. Section 3.1 defines problem formulation and evaluation metrics, including accuracy and diff-bias scores. Section 3.2 presents our proposed Multi-Persona Thinking (MPT) framework.
@@ 3.1 Problem Formulation
In this work, we adopt the widely used problem formulation for LLM debiasing (Nadeem et al., 2021; Parrish et al., 2022; Shaikh et al., 2023; Yang et al., 2025). Each instance consists of a context $c$ and a question $q$, involving two contrasting groups (the target group and the counter-target group). Three answer choices $\{a_{0},a_{1},a_{2}\}$ are given, referring to a target-biased option, a counter-target-biased option, and an unknown option. Example 1 in Table 1 describes a scenario related to age stereotypes, where the correct answer should be $a_{2}$ (unknown) as other options cannot be inferred from the context. This is referred to as an ambiguous setup. The context can be disambiguated, where $c$ contains additional information that explicitly supports one of the answers. In Example 2, the answer clearly becomes “The grandson” with the additional context information.
To measure bias in LLMs, we use the accuracy and diff-bias score (Jin et al., 2024) based on the notation shown in Table 2.
@@ 3.2 Multi-Persona Thinking Framework
We propose the MPT framework based on the assumption that a single LLM encodes a range of social perspectives and associated biases inherently  Gupta et al. (2024). MPT provides a procedure to allow the internal perspectives to exchange viewpoints and reconcile bias during the model’s own reasoning process. We first instantiate different personas with specific system prompts. Then we instruct the model to reason from these perspectives as a simulated self-debate. Finally, the model reviews these viewpoints and provides a final self-debiased answer. Since this is an inference-time framework, it is lightweight and broadly applicable. Details of our approach are introduced below.
@@ 3.2.1 Persona Initialization
For each input example, we define three personas, $P=\{p_{1},p_{2},p_{3}\}$. Two personas correspond to the contrasting social groups specified in the dataset metadata (examples are shown in the Appendix A, Table 9), and an additional neutral general public persona. This design creates a controlled contrast Galinsky and Moskowitz (2000). The two group-specific personas allow the model to surface potential latent biases from opposite perspectives, whereas the neutral persona helps prevent the reasoning process from being dominated by either side. As a result, the final decision is based on a more balanced view.
In our framework, each persona $p_{i}$ is assigned through a system prompt. Examples of the prompts are provided in the Appendix E.
@@ 3.2.2 Iterative Reasoning
At the beginning ($t=0$), the model independently adopts each persona $p_{i}$ to generate an initial textual response $r_{i}^{(0)}$ and an answer $y_{i}^{(0)}\in\{a_{0},a_{1},a_{2}\}$ based on the question prompt $P_{\text{question}}^{(0)}$:
@@
Following the initial generation, the process proceeds through $R$ rounds of dialectical reasoning: In each subsequent round ($t>0$), the model adopting persona $p_{i}$ is instructed to generate a refined response $r_{i}^{(t)}$ and answer $y_{i}^{(t)}$, given a review prompt $P_{\text{review}}^{(t)}$ containing responses from all personas in the previous round $\{r_{j}^{(t-1)}\}_{j\in\{1,2,3\}}$:
@@
The initialization aims to capture the instinctive and potentially biased answer from each perspective. This makes the model’s internal stereotypes and bias explicit and observable. The subsequent $R$ rounds simulate a cognitive debiasing (Soll et al., 2015): the model is instructed to engage in a self-correction by reasoning from different personas to review and confront the contrasting arguments.
@@ 3.2.3 Final Aggregation
In the final aggregation step, the model is no longer constrained by explicit identities, and it acts as a neutral “judge”. It reviews the final arguments from all the personas and integrates them into one final answer. This encourages the model to select the most logical and fact-based conclusion, rather than simple majority voting, which improves both the quality and the objectivity of the output.
@@ 4.1 Datasets
We perform our evaluation on two widely used stereotyping benchmarks: BBQ (Parrish et al., 2022) and StereoSet (Nadeem et al., 2021).
@@
StereoSet is a benchmark comprising 16,995 crowdsourced context association tests that measures language modeling ability and stereotype bias across four domains: gender, profession, race, and religion. The benchmark includes two task formats: word-level (fill-in-the-blank) and sentence-level (choose the most logical continuation). In its original setup, each instance provides a context with three candidate associations: one stereotypical, one counter-stereotypical, and one unrelated. To make our evaluation focus on bias measurement rather than general language modeling, we adapt the test set by replacing the “unrelated” option with “unknown”. More details on the datasets are provided in the Appendix A.
@@ 4.2 Baselines
We compare MPT with representative baselines categorized by their underlying prompting strategies.
Direct prompting. The model is prompted in a single step without iterative refinement. We consider three variants to assess the effectiveness of single-turn prompting: (i) Standard prompting, which provides only the task description. (ii) Explicit debiasing, which adds explicit instructions to avoid bias. (iii) Persona-based prompting, which instructs the model to avoid bias by parallel thinking from the perspectives of the two contrasting social groups.
Self-consistency (SC; Wang et al., 2022). This approach samples multiple independent outputs from direct prompting under stochastic decoding. The final answer is determined by majority vote. We apply self-consistency to all three variants of direct prompting.
Re-prompting (Gallegos et al., 2025). Re-prompting is a two-stage prompting strategy where the model first produces an initial answer, and then is re-invoked with additional explicit debiasing or persona-based instructions to reconsider and potentially revise its response.
Multi-agent debate (MAD; Du et al., 2023). MAD is a framework in which multiple model instances engage in debate and refine their reasoning through iterative exchanges. Unlike our method, this baseline does not explicitly model diverse social perspectives for bias mitigation.
All prompts and templates are detailed in the Appendix E.